OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-3, -1) with a slope of 4. y = 4x - 11 y = 4x + 11 y = 4x + 7 y = 4x - 7

OpenStudy (mathstudent55):

The equation of a line in slope-intercept form is y = mx + b, where m = slope and b = y-intercept. Since you know that the slope is 4, you have y = 4x + b Plug in (-3, -1) for x and y, and solve for b. Then compare it with the choices.

OpenStudy (anonymous):

-1 = 4(-3) + b what do i do after

OpenStudy (mathstudent55):

Multiply the 4 and -3 on the right side.

OpenStudy (anonymous):

-12

OpenStudy (mathstudent55):

So you have -1 = -12 + b Now add 12 to both sides to get rid of -12 from the right side.

OpenStudy (anonymous):

11 = b

OpenStudy (mathstudent55):

Right. Since the equation is y = 4x + b, and b = 11, the equation is y = 4x + 11

OpenStudy (anonymous):

ohhh i get it! thank you so much!

OpenStudy (mathstudent55):

You're welcome. Good job!